HI I AM AMRIT OF 7TH CLASS FROM INDIA
MY QUESTION IS
THE PERIMETER OF A RECTANGLE IS 100 CM AND LENGTH IS 40 CM FIND BREADTH

Answer by Theo(13342)

perimeter = 2 times length + 2 times width.
width and breadth mean the same thing, i believe.

perimeter is equal to 100
length is equal to 40

formula becomes:

100 = 2 * 40 + 2 * w

w represents the width.

solve for w as follows:

original equation:
100 = 2*40 + 2*w
simplify to get:
100 = 80 + 2*w
subtract 80 from both sides of the equation to get:
20 = 2*w
divide both sides of the equation by 2 to get:
10 = w which is the same as:
w = 10

the length is 40 and the width is 10.

2 * 40 = 80
2 * 10 = 20
80 + 20 = 100

your answer is that the breadth is equal to 10 cm.